# A Statistical Physics perspective on fairness in shared expenses: The bar bill analogy

### v1

@@ 2 Statistical model and Thermodynamic analogy
[…]
@@ 3 Canonical analysis of statistical quantities
Following the definition of Eq. (4), the partition function related to the function (7) is given by
@@
The entropy can also be calculated analytically. By analogy with the canonical ensemble [3], we obtain $S_{\alpha}=\alpha\,\langle x\rangle_{\alpha}+\ln Z_{\alpha}$  [^2] Although discussed here in the context of the gamma distribution, this expression holds exactly for any underlying distribution $\rho(x)$. In cases where analytical evaluation is not feasible, it can still be computed numerically via sampling of $\rho(x)$ combined with the Boltzmann weighting $e^{-\alpha\,x}$.. Considering Eqs. (8) and (9), one obtains
@@ 4 Simulation results
In Fig. 3 we plot $\delta_{N}$ vs $N$ for typical values of $\alpha$. As it can be seen, our numerical results confirm that, regardless of the value of $\alpha$, the relative deviation consistently scales as $\delta_{N}\sim N^{-a}$, with the exponent $a$ very close to 1/21/21 / 2. This finding is in agreement with the central limit theorem and indicates that, even under canonical weighting by the Boltzmann factor $e^{-\alpha\,x}$, individual expenditures remain effectively independent. The system thus retains the statistical extensivity of an ideal gas, where aggregate fluctuations are suppressed as the number of individuals increases [3]. Thus, as $N$ increases, deviations shrink and fairness in the bill improves.
Figure 3: Relative deviation $\delta_{N}$, given by Eq. (16), vs group size $N$, for typical values of $\alpha$, plotted in the log-log scale. As a guide to the eye, a straight line with slope 0.5-0.5- 0.5 is shown. Each point represents an average over $10^{6}$ independent simulations.
[…]
@@
$00$ 20.00\\$\,20.00\$ 20.00 \$ 14.14\\$\,14.14\$ 14.14 \$ 4.47\\$\,4.47\$ 4.47 \$ 2.58\\$\,2.58\$ 2.58
0.050.050.05 \$ 13.33\\$\,13.33\$ 13.33 \$ 9.43\\$\,9.43\$ 9.43 \$ 2.98\\$\,2.98\$ 2.98 \$ 1.72\\$\,1.72\$ 1.72
0.100.100.10 \$ 10.00\\$\,10.00\$ 10.00 \$ 7.07\\$\,7.07\$ 7.07 \$ 2.24\\$\,2.24\$ 2.24 \$ 1.29\\$\,1.29\$ 1.29
0.200.200.20 \$ 6.67\\$\,6.67\$ 6.67 \$ 4.71\\$\,4.71\$ 4.71 \$ 1.49\\$\,1.49\$ 1.49 \$ 0.86\\$\,0.86\$ 0.86
0.300.300.30 \$ 5.00\\$\,5.00\$ 5.00 \$ 3.54\\$\,3.54\$ 3.54 \$ 1.12\\$\,1.12\$ 1.12 \$ 0.65\\$\,0.65\$ 0.65
@@ 5 Discussion and conclusions
To sum up the previous sections, in this work we considered a model based on an analogy with the ideal gas in the canonical ensemble. In this framework, the consumption $x_{i}$ of a given individual is analogous to the kinetic energy of a gas particle $E_{i}$, and the average consumption corresponds to the average energy per particle. A Boltzmann-like factor is introduced, namely $e^{-\alpha\,x_{i}}$, where the parameter $\alpha$ can be interpreted as a penalization factor. This parameter plays a role which is analogous to the inverse temperature $\beta=1/T$ in statistical mechanics. At low $\alpha$ (high temperature), individual expenditures are weakly constrained and the system exhibits large fluctuations, akin to a high-temperature gas where particles have a broad energy distribution. Conversely, at high $\alpha$ (low temperature), the Boltzmann weighting strongly penalizes large expenditures, inducing a more concentrated, “ordered” distribution of behavior, analogous to how particles in a low-temperature gas predominantly occupy lower-energy states. We considered the gamma distribution for the density of states, where an state in the model is given by the expenditures of the $N$ individuals in the group, i.e., $\{x\}=\{x_{1},x_{2},...,x_{N}\}$.
@@
Numerical simulations for various values of the penalization parameter $\alpha$ confirm that the relative deviation of the group-averaged bill $\delta_{N}$ of the canonical average scales with the group size $N$ as $\delta\sim N^{-a}$, with the exponent $a$ being very close to 1/21/21 / 2. This indicates that statistical independence among individual expenditures is preserved even under canonical weighting, consistent with the law of large numbers [18]. Such behavior highlights the robustness of the collective average against individual variability, further supporting the analogy with extensive thermodynamic systems [3]. In fact, the social aspect of a bar bill interaction, e.g. noticing if one’s consumption differs too much from the others may hinder this ones next orders, may contribute to regulate the system as a whole and keep it inside the low temperature regime (i.e. high $\alpha$) of the canonical ensemble.
@@
The apparent fairness that arises when a group divides a bill equally can be seen as a signature of self-organization and emergent order, two hallmarks of complex systems. Even though individuals make autonomous, diverse choices about what to consume, the collective outcome displays regularity, predictability, and a sense of equity. This is not imposed externally, but emerges from the interplay of statistical laws, weak social correlations, and ensemble behavior. This regularity is not accidental: it is a hallmark of complex systems where microscopic randomness coupled to self-organizing mechanisms, such as a sense of guilt in spending much more than others, gives rise to macroscopic predictability. Thus, what seems like a simple social custom is in fact rooted in the same principles that govern energy distributions in matter.
@@
In fact, agents operate based on local rules or preferences, yet the system as a whole exhibits coherence and structure [6]. The emergence of fairness from randomness mirrors other examples in nature and society, such as traffic flows, consensus formation, and financial market dynamics. In our model, equal division of the total cost becomes a reliable and robust rule, not because everyone consumes the same, but because the aggregate behavior smooths out individual variation. This insight reinforces the value of statistical physics as a conceptual lens for understanding not just physical matter, but the organized patterns of social and economic life.
Perhaps most strikingly, this study exemplifies the explanatory power of statistical physics in interdisciplinary contexts. The simple act of splitting a restaurant bill among friends can be reinterpreted as a thermodynamic system in statistical equilibrium under social constraints. What seems like an informal social convention – dividing the total cost equally – emerges naturally from deep statistical principles: mutual compensation, suppression of fluctuations, and ensemble averaging. This highlights the remarkable ability of statistical physics to uncover hidden order in everyday human behavior and to provide a unified framework for understanding fairness, cooperation, and stability across seemingly unrelated domains. This may be a stepping stone in understanding, through thermodynamical statistical physics, the emergence of collective behavior in large econophysical systems, as well as shedding a light into daily simple social behaviors.
@@ Acknowledgments
N. Crokidakis acknowledges partial financial support from the Brazilian scientific funding agency Conselho Nacional de Desenvolvimento Científico e Tecnológico, Brazil (CNPq, Grant 308643/2023-2). L. Sigaud acknowledges partial financial support from the Brazilian scientific funding institute INCT-FNA.

### v2

@@ 2 Statistical model and Thermodynamic analogy
Partition function $Z_{\alpha}=\int e^{-\alpha x}\rho(x)\,dx$ $Z_{\beta}=\int e^{-\beta E}g(E)\,dE$
@@
3 The maximum relative entropy method: A natural way of explanation
In the previous section we defined our “penalized” probability for consumption in the form given by Eq. (3), with the corresponding partition function given by Eq. (4).
However, such two equations can be obtained from the maximum relative entropy method, which make the motivation for the proposed model even clearer. Historically, the maximum relative entropy method is a direct descendant of the MaxEnt method, pioneered by Jaynes [16, 17, 18, 19].
We can start from a prior model $\rho(x)$ and update it by maximizing the relative entropy functional [19, 20]
$\mathcal{S}[P,\rho]=-\int_{0}^{\infty}P(x)\,\ln\frac{P(x)}{\rho(x)}\,dx$ (7)
The probability $P(x)$ is subject to the constraint
$\langle x\rangle=\int_{0}^{\infty}x\,P(x)\,dx=x_{0}$ (8)
where $x_{0}$ is the the known value of the average individual consumption. Maximizing $\mathcal{S}[P,\rho]$, Eq. (7), over $P(x)$ subject to (8) and normalization, yields [19]
$P(x)=\frac{1}{\int_{0}^{\infty}e^{-\lambda x}\,\rho(x)\,dx}\,e^{-\lambda x}\,\rho(x)$ (9)
where $\lambda$ is a Lagrange multiplier. This distribution is normalized, and we have the partition function $Z$ as the normalization factor,
$Z=\int_{0}^{\infty}e^{-\lambda x}\rho(x)\,dx$ (10)
Thus, our penalization parameter $\alpha$ emerges as the Lagrange multiplier $\lambda$ imposing the constraint in (8), so its magnitude $|\alpha|$ in fact measures the strength of said constraint, in complete agreement with our initial intuition.
In addition, from the functional (7), one can obtain the entropy $S$,
$S=\left\langle-\ln\frac{P(x)}{\rho(x)}\right\rangle$ (11)
which leads immediately to $S\leq 0$, with $S=0$ occuring only for $P(x)=\rho(x)$, i.e., for $\alpha=0$ (no penalization). We will discuss the meaning of the negative signal of entropy in the next section.
@@ 4 Canonical analysis of statistical quantities
Following Eq. (4), the partition function related to the function (12) is given by
@@
The entropy can also be calculated analytically. From Eq. (3), one obtains $P(x)/\rho(x)=(1/Z)\,e^{-\alpha x}$, which leads to $\ln(P(x)/\rho(x))=-\alpha\,x-\ln Z$. Considering such result, the entropy can be obtained from Eq. (11), leading to $S_{\alpha}=\alpha\,\langle x\rangle_{\alpha}+\ln Z_{\alpha}$  [^2] Although discussed here in the context of the gamma distribution, this expression holds exactly for any underlying distribution $\rho(x)$. In cases where analytical evaluation is not feasible, it can still be computed numerically via sampling of $\rho(x)$ combined with the Boltzmann weighting $e^{-\alpha\,x}$.. Considering Eqs. (13) and (14), one obtains
@@ 5 Simulation results
In Fig. 3 we plot $\delta_{N}$ vs $N$ for typical values of $\alpha$. As it can be seen, our numerical results confirm that, regardless of the value of $\alpha$, the relative deviation consistently scales as $\delta_{N}\sim N^{-a}$, with the exponent $a$ very close to $1/2$. This finding is in agreement with the central limit theorem and indicates that, even under canonical weighting by the Boltzmann factor $e^{-\alpha\,x}$, individual expenditures remain effectively independent. The system thus retains the statistical extensivity of an ideal gas, where aggregate fluctuations are suppressed as the number of individuals increases [3]. Thus, as $N$ increases, deviations shrink and fairness in the bill improves.
Figure 3: Relative deviation $\delta_{N}$, given by Eq. (21), vs group size $N$, for typical values of $\alpha$, plotted in the log-log scale. As a guide to the eye, a straight line with slope 0.5-0.5 is shown. Each point represents an average over $10^{6}$ independent simulations.
As a numerical example, let us consider the same parameters for the gamma distribution adopted throughout the text, i.e., $k=2.0$ and $\theta=10.0$. First, for the unpenalized case ($\alpha=0.0$), the average consumption is given by Eq. (14) with $\alpha=0.0$, yielding $\langle x\rangle_{\alpha=0.0}=k\,\theta=\mathdollar\,20.00$, with standard deviation given by Eq. (15), namely $\sigma_{\alpha=0.0}=\sqrt{k\,\theta^{2}}=\mathdollar\,14.14$. Introducing a moderate penalization ($\alpha=0.05$) reduces the average to \$ $\mathdollar\,13.33$ and the standard deviation to \$ $\mathdollar\,9.43$. Considering a stronger penalization such as $\alpha=0.30$ further decreases the average to \$ $\mathdollar\,5.00$ with a standard deviation of only \$ $\mathdollar\,3.54$. For a group of size $N=10$, for example, the fluctuation of the group-averaged bill, given by $\sigma_{\bar{x}_{N}}=\sqrt{\sigma^{2}_{\alpha}/N}$, reduces accordingly from \$ $\mathdollar\,4.47$ in the unpenalized case ($\alpha=0.0$) to \$ $\mathdollar\,2.98$ at $\alpha=0.05$ and \$ $\mathdollar\,1.12$ at $\alpha=0.30$. This illustrates how penalization leads to both lower average expenditures and increased collective homogeneity. Moreover, as previously discussed, the fluctuation of the group-averaged bill decreases with group size, reinforcing the fairness of each individual paying the group’s average consumption. A summary of these results is presented in Table 2, where we also include the group mean standard deviation for $N=30$ to enable comparison with the results for group size $N=10$. It can be seen that increasing the group size further reduces fluctuations in the group-averaged bill, thereby reinforcing the statistical fairness of equal cost-sharing.
@@
$0.00$ \$ $\mathdollar\,20.00$ \$ $\mathdollar\,14.14$ \$ $\mathdollar\,4.47$ \$ $\mathdollar\,2.58$
$0.05$ \$ $\mathdollar\,13.33$ \$ $\mathdollar\,9.43$ \$ $\mathdollar\,2.98$ \$ $\mathdollar\,1.72$
$0.10$ \$ $\mathdollar\,10.00$ \$ $\mathdollar\,7.07$ \$ $\mathdollar\,2.24$ \$ $\mathdollar\,1.29$
$0.20$ \$ $\mathdollar\,6.67$ \$ $\mathdollar\,4.71$ \$ $\mathdollar\,1.49$ \$ $\mathdollar\,0.86$
$0.30$ \$ $\mathdollar\,5.00$ \$ $\mathdollar\,3.54$ \$ $\mathdollar\,1.12$ \$ $\mathdollar\,0.65$
@@
These results are corroborated by previous researches into social behavior that consider the bill-splitting scenario. In fact, many studies point to the fact that in groups, specially large ones, food and drink consumers tend to flock to particular categories of items (even if choosing different items themselves) [23], which can lead to discrepancies that average out when the whole bill is divided [24]. This was what was observed in studies made only with receipts, but in more careful probing using the examination of who ordered which item first, it was further observed that the leading orders tend to influence the other consumers in the group in their respective choices [25, 26]. In fact, both the social experience and the social ties between the consumers seem to be enhanced by the bill splitting [24].
On the other hand, it was also shown very recently that the expected happiness of paying a bill in full for a cherished person is even more rewarding [27]. Nevertheless, when grouped with someone with weaker social ties, the splitting of the bill can still be the preferred option. Furthermore, a study by Y. Shani [28] corroborates the implications proposed by Gneezy et al. [2], but in a controlled situation where people are questioned solely about buying chocolate candy bars, which does not reproduce the informality and spontaneity of a dining environment. Still, one must take into account that what the results here show is a tendency of the discrepancies between different consumptions to even out when splitting a bill equally - the sharing experience and the influence of previous orders [23, 24, 25, 26] are maybe natural social causes of that result, which in turn are modeled here by a thermodynamical system. Of course, this differs considerably from both the works of Gneezy et al. [2] and Shani [28] and from market econophysical models, where the change from a social to a market-influenced setting leads to a competitive behavior for which our model is not intended [29].
@@ 6 Discussion and conclusions
To sum up the previous sections, in this work we considered a model based on an analogy with the ideal gas in the canonical ensemble. In this framework, the consumption $x_{i}$ of a given individual is analogous to the kinetic energy of a gas particle $E_{i}$, and the average consumption corresponds to the average energy per particle. A Boltzmann-like factor is introduced, namely $e^{-\alpha\,x_{i}}$, where the parameter $\alpha$ can be interpreted as a penalization factor. Taking into account a functional entropy form, and the maximum relative entropy method, the parameter $\alpha$ was identified as the Lagrange multiplier used in the maximization procedure of the functional over the probability $P(x_{i})$, subject to normalization and the constraint given by the average consumption. The parameter $\alpha$ plays a role which is analogous to the inverse temperature $\beta=1/T$ in statistical mechanics. At low $\alpha$ (high temperature), individual expenditures are weakly constrained and the system exhibits large fluctuations, akin to a high-temperature gas where particles have a broad energy distribution. Conversely, at high $\alpha$ (low temperature), the Boltzmann weighting strongly penalizes large expenditures, inducing a more concentrated, “ordered” distribution of behavior, analogous to how particles in a low-temperature gas predominantly occupy lower-energy states. We considered the gamma distribution for the density of states, where an state in the model is given by the expenditures of the $N$ individuals in the group, i.e., $\{x\}=\{x_{1},x_{2},...,x_{N}\}$.
@@
Numerical simulations for various values of the penalization parameter $\alpha$ confirm that the relative deviation of the group-averaged bill $\delta_{N}$ of the canonical average scales with the group size $N$ as $\delta\sim N^{-a}$, with the exponent $a$ being very close to $1/2$. This indicates that statistical independence among individual expenditures is preserved even under canonical weighting, consistent with the law of large numbers [30]. Such behavior highlights the robustness of the collective average against individual variability, further supporting the analogy with extensive thermodynamic systems [3]. In fact, the social aspect of a bar bill interaction, e.g. noticing if one’s consumption differs too much from the others may hinder this ones next orders, may contribute to regulate the system as a whole and keep it inside the low temperature regime (i.e. high $\alpha$) of the canonical ensemble.
@@
Thus, even though individuals make autonomous, diverse choices about what to consume, the collective outcome displays regularity, predictability, and a sense of equity. This is not imposed externally, but emerges from the interplay of statistical laws, weak social correlations, and ensemble behavior. This regularity is not accidental: it is a resemblance to complex systems where microscopic randomness coupled to self-organizing (in this case, social) mechanisms, such as a sense of guilt in spending much more than others, gives rise to macroscopic predictability. Thus, what seems like a simple social custom is in fact rooted in the same principles that govern energy distributions in matter.
@@
In fact, agents operate based on local rules or preferences, yet the system as a whole exhibits coherence and structure [6]. The emergence of fairness from randomness mirrors other examples in nature and society, such as traffic flows, consensus formation, and opinion dynamics. In our model, equal division of the total cost becomes a reliable and robust rule, not because everyone consumes the same, but because the aggregate behavior smooths out individual variation. This insight reinforces the value of statistical physics as a conceptual lens for understanding not just physical matter, but the organized patterns of social and economic life. The simple act of splitting a restaurant bill among friends can be reinterpreted as a thermodynamic system in statistical equilibrium under social constraints. What seems like an informal social convention – dividing the total cost equally – emerges naturally from simple statistical principles: mutual compensation, suppression of fluctuations, and ensemble averaging.
@@ Acknowledgments
N. Crokidakis acknowledges partial financial support from the Brazilian scientific funding agency Conselho Nacional de Desenvolvimento Científico e Tecnológico, Brazil (CNPq, Grant 308643/2023-2). L. Sigaud acknowledges partial financial support from the Brazilian scientific funding institute INCT-FNA. We are indebted to an anonymous referee who drew our attention to the application of the maximum relative entropy method to our model.